can you please explain to me how this problem can be solved: a square has a 40 cm diagonal .How long is each side of the square ? Round your answer to the nearest of a centimeter .
Answer by checkley77(12844)

THE 40 CM IS THE HYPOTENUSE [C] OF A RIGHT TRIANGLE.
A=B (THE 2 SIDES ARE EQUAL).
A^2+B^2=C^2
2A*2=40
A^2=40/2
A^2=20
S=SQRT20
A=4.47 OR 4 ANS.